### v1

# Algorithmic Misjudgement in Google Search Results: Evidence from Auditing the US Online Electoral Information Environment

Abstract.
Google Search is an important way that people seek information about politics (Dutton et al., 2017), and Google states that it is “committed to providing timely and authoritative information on Google Search to help voters understand, navigate, and participate in democratic processes”. [^1] https://elections.google/civics-in-search/ In this paper, we interrogate the extent to which government-maintained web domains are represented in the online environment of electoral information of the 2022 US midterm elections, as captured through Google Search results in 3.45 million SERPs for 786 locations across the United States between October and November 2022. Although we find that almost 40% of organic results are contributed by the 40% of government domains, this proportional equilibrium hides the fact that most results either belong to a small number of popular domains or are mistargeted (at a rate of 71.18%) with respect to the location of the search. We consider the frequent omission and mistargeting of non-federal websites a form of algorithmic misjudgement that contributes to civic harm, by obscuring the important role that these institutions play in the election information environment.
algorithm auditing, problematic behavior, google search audit, elections, algorithmic misjudgement
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ACM Conference on Fairness, Accountability, and Transparency; June 03–06, 2024; Rio de Janeiro, Brazilisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Electoral denialism (the belief that elections are rigged or that electoral fraud is rampant) in the United States (but also other countries, such as Brasil and Peru) is emerging as a significant threat to democracy (Cameron, 2023), and it has roots, among others in racial resentment and rising conspiracism (Stewart III, 2023). Since the early 2000s, internet was regarded as a medium “uniquely receptive to conspiracism” (James, 2000), but in the past two decades, there is little doubt that it has contributed significantly to conspiracism going mainstream (Rosenblum and Muirhead, 2019). Following the 2020 US Presidential Election, “election rigging” conspiracies that came to be labeled as “The Big Lie” led among others to the violent attack to the US Capitol on January 6 2021 (Fahey, 2023). While beliefs about electoral fraud are somewhat persistent (Enders et al., 2021), the 2020 electoral denialism movement was different because it was embraced by partisan state legislatures, as reported by the Brennan Center for Justice. [^2] https://www.brennancenter.org/our-work/research-reports/voting-laws-roundup-june-2023 As a result, a large number of changes to electoral laws (some of them restricting and some of them expanding voting rights), became operational ahead of the 2022 US midterm election. [^3] https://www.reuters.com/legal/how-new-us-laws-could-trip-up-voters-midterm-elections-2022-11-01/
In an environment of election denialism on one hand and changing electoral laws on the other, voters must seek accurate, up-to-date, and relevant information about elections and voting. A survey by the Bipartisan Policy Center in Fall 2022 found that voters are “most likely to look to their state and local election officials, and search engines” for election information, [^4] https://bipartisanpolicy.org/blog/new-survey-data-election-information/ aligning with previous research (Dutton et al., 2017). Election officials similarly position themselves as authorities on election information. For example, the National Association of Secretaries of State launched a campaign ahead of the 2020 presidential elections called #TrustedInfo2020 to “promote election officials as the trusted sources of election information.” Operating as #TrustedInfo during non-federal election years and now #TrustedInfo2024 ahead of the 2024 presidential elections, NASS states, “By driving voters directly to election officials’ websites, social media pages, and materials, they will be able to receive credible, timely information on each step of the elections process.” [^5] https://www.nass.org/initiatives/trustedinfo
Likewise, search engines also position themselves as trustworthy intermediaries for seeking election information. Google states that it is “committed to providing timely and authoritative information on Google Search to help voters understand, navigate, and participate in democratic processes.” [^6] https://elections.google/civics-in-search/ In an article about their work to “support” the 2022 US midterm elections, Google’s Vice President for Trust and Safety writes, “Our work is centered around connecting voters to the latest election information …” [^7] https://blog.google/outreach-initiatives/civics/our-ongoing-work-to-support-the-2022-us-midterm-elections/ and highlights “features that show data from nonpartisan organizations” that work to “connect voters with accurate information about voter registration and how to vote,” referring to Google Search election information panels that contain aggregated information from Democracy Works [^8] Democracy Works describes itself as “a nonpartisan, nonprofit organization that collaborates with election officials, leading tech platforms, and world-class partners to drive voter access and participation”. Source: https://www.democracy.works/about. about where and how to vote.
Study Motivation: Given this background, three assumptions motivate the current study: (1) voting-eligible US citizens who search about electoral information using Google’s search engine have the right to receive accurate and relevant search results, regardless of where they live, 2) similar to Ballatore et al. (Andrea Ballatore and Sen, 2017), we acknowledge a spatial dimension to searching and search results, in which users searching for information related to elections and voting deserve results relevant to their geographic area, and 3) given the size of the United States and the varying voting laws per state (most of them updated ahead of the 2022 elections), receiving information from authoritative, locality-specific sources is a reasonable expectation. With these in mind, we proceed to audit Google Search results to determine the extent to which government-maintained web domains, as authoritative sources on election information, are represented in the online environment of electoral information, during the 2022 US midterm elections. Specifically, we analyze where, how often, and how accurately state, county, and local government domains appear in Google’s organic results for election-related queries collected in 786 locations across the United States between October and November 2022, in a dataset of 3.45 million SERPs. Drawing parallels to studies of geotargeting errors in advertisements (Bandy and Hecht, 2021), we label the misalignment between the geographic area that a non-federal government web domain serves (such as a state or county) and the location for which it appears in search results as an instance of mistargeting, in which the search algorithm targets locality-specific information to a user in a different (incorrect) location. In this paper, we answer two related research questions:
  • RQ1: To what extent are government web domains represented in the online environment of electoral information about the 2022 US Midterm Elections, as mediated by Google Search?
  • RQ2: To what extent are state, county, and local government web domains correctly targeted to their respective constituents?
Findings and Contributions: We find that 40% of domains that appear among organic results are government websites, constituting nearly 40% of all organic results. However, in the 32.89% of organic results contributed by non-federal government domains, there is a 71.18% mistargeting rate. Our detailed analysis indicates a lack of consistency in the search algorithm’s treatment of locality-specific government sources, leaving most government domains, especially at the local level, appearing minimally, inconsistently, or missing entirely from election-related search results. We frame the omission and frequent mistargeting of non-federal government domains as a form of algorithmic misjudgement that obscures the important role that these domains play in providing official, locally-relevant information to voters. To the best of our knowledge, this is the largest locality-specific Google Search audit in the context of a political election.
2. Literature Review
2.1. Algorithm Auditing and Problematic Behavior
Online platforms and other sociotechnical and algorithmic systems are neither neutral nor unbiased entities. As Gillespie (Gillespie, 2010) argues, the presentation of platforms as open or neutral parties obscures the consequences of deliberate choices platforms make as to what content is published and how content is shown. Through a systematic literature review, Bandy organizes potential problematic behavior of algorithmic systems into four main categories (a) discrimination based on identity or socioeconomic status; (b) distortion or obscuration of reality, (c) exploitation of sensitive user information, and (d) misjudgement and inaccurate classifications (Bandy, 2021). Potential consequences of algorithmic systems become more wide-reaching when considering the interplay between the technical components of algorithmic systems and the social systems in which they exist; one such theme of sociotechnical harm is social systems harms, including how algorithmic systems contribute to the spread of misinformation, disinformation, and malinformation (informational harms) and how algorithmic systems impact governance, democracy, and civil liberties (political and civic harms) (Shelby et al., 2023).
Sandvig et al. (Sandvig et al., 2014) propose algorithm auditing as a way of studying problematic behavior in Internet platforms and algorithmic systems, particularly problematic behavior that is subtle or unobservable via one-off instances. To date, search engines are among the most audited algorithmic systems, often for discrimination or distortion problematic behavior. Prior research has focused on personalization of search results, including personalization based on logged-in status (Hannak et al., 2013), browsing history (Le et al., 2019; Haim et al., 2018), and location (Kliman-Silver et al., 2015); composition of search results, including the presence and content of snippets (Hu et al., 2019), Top stories (Lurie and Mustafaraj, [n. d.]; Trielli and Diakopoulos, 2019), and other non-organic panels (Robertson et al., 2018b); and prominence of different types of sources among organic results (Fischer et al., 2020; Courtois et al., 2018; Metaxa et al., 2019). The pervasive concern in such studies is that the algorithmic choices are leading to a distorted view of reality. Meanwhile, studied to a lesser extent is problematic behavior related to algorithmic misjudgement. As Bandy defines it (Bandy, 2021), misjudgement occurs when an algorithm makes incorrect predictions or classifications, and misjudgement often leads to other types of problematic behavior. Algorithmic misjudgement has been studied in the criminal justice context (eg. recidivism prediction (Duwe and Kim, 2017)), as well as the advertising context, where algorithms incorrectly infer information about users (Bandy, 2021). One implication of advertising algorithm misjudgement is that purchased advertisements do not reach intended audiences. For example, Bandy and Hecht (Bandy and Hecht, 2021) find that geotargeting errors are common in Google Display Network, with users often seeing advertisements not intended for their ZIP code. However, they find that severity of such geotargeting errors were minor in the context of advertising goals, with most users living in the intended county, designated market area, or state (Bandy and Hecht, 2021). With this study, we contribute to the literature on algorithmic misjudgement in the context of location-sensitive electoral information.
2.2. Politics and Search Engines
Much of previous work on auditing search engines is rooted in a political context, motivated by the potential informational, political, and civic harms that could result from biased search. Epstein and Robertson (Epstein and Robertson, 2015) hypothesized, for example, that biased search results can impact electoral outcomes by shifting voter preferences based on manipulated rankings, as studied in controlled, lab-based experiments (Epstein and Robertson, 2015; Epstein et al., 2017). However, algorithm audits in the wild find little evidence for this concern (Robertson et al., 2018a, b). Findings that support the presence of filter bubbles in Google Search and Google News are also limited (Robertson et al., 2018a; Puschmann, 2019; Courtois et al., 2018; Haim et al., 2018; Trielli and Diakopoulos, 2022). Trielli and Diakopoulos (Trielli and Diakopoulos, 2022) in particular find that while members of different ideological groups use different queries to search for political information, such partisan queries do not lead to a filter bubble effect; rather, they observe a “mainstreaming effect”, where highly similar results appear even for diverging partisan queries. Moreover, Robertson et al. (Robertson et al., 2023) find that user choice plays a larger role in engagement with partisan or unreliable news than algorithmic curation, further evidence against the idea of partisan filter bubbles.
Given the focus on potential bias or filter bubbles, the query phrases used in most political information audits are intended to capture such phenomena with respect to either the candidates or issues central to an election. Examples include queries related to Donald Trump’s inauguration (Robertson et al., 2018b), candidate names (Metaxa et al., 2019), partisan terms about immigration (Le et al., 2019), socio-political themes, (Courtois et al., 2018), and crowd-sourced queries from users about senate candidates (Trielli and Diakopoulos, 2022). Most of the audits that focus on candidate names are conducted during election periods, including the 2016 US Congressional Elections (Metaxas and Pruksachatkun, 2017), the 2018 US Midterms (Trielli and Diakopoulos, 2022; Metaxa et al., 2019), and 2020 US Presidential Primaries (Urman et al., 2022; Kawakami et al., 2020). Diakopoulos et al. (Diakopoulos et al., 2018) and Lurie and Mulligan (Lurie and Mulligan, 2021) similarly study various information panels that appear in Google search results when searching for candidate or representative names, with the latter finding that featured snippets are often likely to mislead searchers by presenting incorrect or incomplete information about a representative. Lurie and Mulligan contextualize this behavior as “algorithmic breakdown”, and identify place name ambiguity as an important contributor to such breakdown, wherein the algorithmic system is unable to discern a particular location (and the respective locality-relevant information, like the representative of a congressional district) from an ambiguous place name, such as a county and city that share the same name. Diakopoulos et al. (Diakopoulos et al., 2018) additionally study the composition of organic results for queries of Democrat and Republican candidates ahead of the 2016 US presidential election, finding that official sources, such as campaign websites, were clustered at the top of organic results in high ranking positions. Most of these studies focus on query phrases selected by researchers in various ways. Mustafaraj et al. (Mustafaraj et al., 2020) discuss that voters have information needs not previously captured in such audits. Building upon prior literature, we create an election-relevant query phrases dataset composed of popular queries from Google Trends and queries formulated by users as reported in (Mustafaraj et al., 2020).
2.3. Source Geoprovenance
As Ballatore et al. (Andrea Ballatore and Sen, 2017) describe, search engines “increasingly mediate not just information but also spatial knowledges and experiences,” as users rely on search engines to inform them about destinations, where to shop, or other ways to interact with their “lived everyday geographies” - including participating in civic processes, such as voting. Moreover, where information originates from, particularly if information originates from local or non-local sources, can impact how people learn about and interact with places near them, to the extent that local community members and non-community members view their locality differently (Sen et al., 2015). The geographic origin of information, or geoprovenance (Sen et al., 2015), has largely been studied in places where volunteered geographic information (e.g. geotagging) is accessible, such as Wikipedia (Sen et al., 2015) and Twitter (Johnson et al., 2016). Ballatore et al. (Andrea Ballatore and Sen, 2017) study the extent to which different country variants of Google Search direct users to locally produced information about places, the first such audit study of search engines. Using a geoprovenance inference algorithm developed by Sen et al. (Sen et al., 2015), they identify the geoprovenance of a URL at the country level and compare URL geoprovenance to the country of search, finding a greater degree of local content for wealthy and well-connected countries. This study uses the concept of geoprovenance of a URL to study algorithmic misjudgement in search results. We audit not only the extent to which government sources are present in search results, but also the extent to which state, county, and local government sources are shown to the geographic areas which they serve.
3. Data Collection
An audit comprises a set of inputs and the corresponding outputs. Our inputs are the query phrases and the geographical locations where Google searches were performed. The outputs are the search engine results pages (SERPs) generated for each query-location pair. We describe how we selected the queries and locations and how we performed the audit.
3.1. Selection of Queries
In line with previous research about centering voters in algorithm audit methodology (Mustafaraj et al., 2020), we developed a list of queries aimed at capturing the diverse information needs of voters who use Google to search for election information.
In early October 2022, we collected Top and Rising related terms from Google Trends for the seed queries “midterm elections”, “voting near me”, and “ballot”, as well as from the Google Trends category “United States election, 2022 - General election”. The location of Google Trends was the United States. From the received results, we excluded terms that were irrelevant to the topic of elections; terms that included the names of political figures or candidates; and any terms that included specific states, state abbreviations, or localities so that queries generalize across locations.
A review of the query list after these steps revealed a lack of general queries about candidates and key voter issues, as well as a lack of biased and/or partisan queries. To fill these gaps, we augmented our query list by manually selecting queries formulated by eligible US voters included in the Voter Searches dataset in (Mustafaraj et al., 2020). Lastly, a few additional seed queries were added related to the topic of abortion, which was one of the top voting issues [^9] https://www.pewresearch.org/politics/2022/08/23/abortion-rises-in-importance-as-a-voting-issue-driven-by-democrats/ in the 2022 election. These processes resulted in a final list of 161 queries, of which 74 were from Google Trends. [^10] The query phrases and their provenance provenance information is available here: https://bit.ly/audit-queries.
Queries were reviewed qualitatively by one author to develop three major themes of a) Voting Logistics (in which queries are primarily about logistical information related to voting); b) Candidates and Issues (in which queries are primarily about candidates, key voting issues, and candidate’s opinions, stances, and/or leanings) and c) General Election Topics, (where queries are primarily general inquiries about the midterm election, such as what midterms are and when they occur). Three labelers then independently labeled all queries into these three categories. The majority label was used as the final label per query. Table 1 shows the query distribution by category and some examples for each category.
Query Category Example Queries n
Voting Logistics voting early near me, sample ballot for my county, what district am i in 68
Candidates and Issues who is on the ballot, candidate midterm info, candidates for my district 50
General Election Topics 2022 midterm, mid term elections, should i vote, 43
Table 1. The audit used 161 queries that cover three major themes related to logistics of voting, candidates and key voting issues, and general inquiries about midterms. The query categories were assigned by three independent labelers.
3.2. Selection of Locations
We selected locations for data collection to correspond to U.S states and congressional districts whose House or Senate midterm races were classified as toss-ups, leaning Democrat, and leaning Republican, according to the Cook Political Report.111111 https://www.cookpolitical.com/ratings/senate-race-ratings and https://www.cookpolitical.com/ratings/house-race-ratings We identified 30 congressional districts with competitive House races across 23 states, as well as 10 states with competitive Senate races; 3 states had only competitive Senate races, making for a total of 26 unique states.
For house races, because congressional district borders do not align with county borders, we first manually compiled a list of counties within the relevant congressional district, according to Wikipedia, [^12] Eg. https://en.wikipedia.org/wiki/California's_27th_congressional_district including only the counties that were described as having most or all of their localities within the congressional district. Next, we identified the associated Federal Information Processing System (FIPS) Codes for each county [^13] https://transition.fcc.gov/oet/info/maps/census/fips/fips.txt and used these codes to filter 2020 U.S. Census data [^14] Dataset titled “Incorporated Places and Minor Civil Divisions Datasets: Subcounty Resident Population Estimates: April 1, 2020 to July 1, 2021 (SUB-EST2021)” from https://www.census.gov/data/tables/time-series/demo/popest/2020s-total-cities-and-towns.html to include localities from the relevant counties only. Additionally, we filtered locations to only include cities, towns, and townships, as identified on the Census.
To sample locations, we identified the median population of possible locations, and selected (up to) 15 locations below the median population and (up to) 15 locations above the median population, as well as the location with the highest population. Some states (for example, Nevada, which had only a competitive Senate race) had fewer than 31 cities, towns, and townships from the relevant counties. This process resulted in 872 locations across 26 states. Google Places API was used to get latitude and longitude coordinates for each of the locations. This information is needed to change the location of the Chrome browser to the desired location for the audit, [^15] Location information is available here: https://bit.ly/audit-locations as shown next.
3.3. Scraping Audit and SERP Validation
We use the Python-scripted Selenium web driver to automate the process of querying Google Search, changing the geolocation, and saving the SERPs. For each query, our custom script opens a blank-slate of the Chrome browser, enters the query, and scrolls to the bottom of the page to update the location using latitude and longitude coordinates (Figure 1). The page reloads, displaying the results for the new location, and the entire HTML page is stored.
Figure 1. The script automatically updates the location with latitude and longitude coordinates so that the search location is “from your device” rather than IP address. Previous audits (Kliman-Silver et al., 2015) have tested and confirmed that search results are personalized based on provided coordinates, rather than the IP address.
SERPs were collected for 30 days, starting on October 18, 2022 and ending November 18, 2022. Data were not collected on November 14 and 15. Data were collected using 12 desktop computers, with 2 to 3 states collecting per computer. Because of the large amount of locations, data collection started each day at 2am EST and continued until the evening. The order of locations was not changed daily, so SERPs per query were collected at approximately the same time within each location, and locations were completed sequentially. 3,838,128 SERPs were collected in total.
To validate that locations were changed correctly, we extracted the text of the location box on each SERP and compared it to the intended location. The validation process revealed two types of errors: (1) the location of a SERP failed to update properly for unknown reasons, and (2) the latitude and longitude coordinates from the Google Places API did not exactly match the actual intended location. To resolve these issues, we remove all 86 locations that had errors and conduct analysis on 3,455,822 SERPs from the remaining 786 locations across 25 states.
4. Methods
Since our research questions revolve around the presence of government-maintained websites in SERPs, we describe in this section how we extracted their occurrences; labeled them according to the government type; and defined when they are mistargeted with respect to the search location.
4.1. Extracting domains and election information boxes
SERPs are composed of multiple components that supplement the organic search results (the blue links) such as top stories, videos, images, the people also ask panel, tweets, and more. During election seasons, Google includes election-specific boxes on many SERPs with information about how to vote, notable dates, or interactive tools to find your voting location. Thus, we parsed SERPs for two types of information: a) the organic search results links; and b) the election information boxes.
Organic search links. 22,750 unique links were extracted from 31,310,263 organic results. Once all such links were extracted from all SERPs, they were cleaned to their root domain; however, given our focus on various government sources, in some cases we maintain the subdomain and avoid aggregating to the root, particularly if the subdomain is a county or town website. For example, while clintonco.illinois.gov is a subdomain of the state government website illinois.gov, we maintain the county government website. 4,556 unique domains were identified.
Election Information Boxes. We identified seven types of election boxes in the SERPs, which were displayed either at the top of the page (with a specific banner) or on the right side. The content of these boxes came from Democracy Works (for notable dates, where to vote, how to vote, etc.), Wikipedia (general information about the election), and Associated Press (election results).
4.2. Identifying Government Websites
To identify government websites, we make use of two repositories maintained by the US government: (1) a list of nearly 10,000 .gov domains maintained by the Cybersecurity and Infrastructure Security Agency (CISA), [^16] https://github.com/cisagov/dotgov-data/blob/main/current-full.csv and (2) a list of 9,200 government-maintained websites outside of the .gov and .mil domains, maintained by the US General Services Administration (GSA). [^17] https://github.com/GSA/govt-urls/blob/main/1_govt_urls_full.csv Both datasets include information about the type of government domain, such as state, county, or local, the agency that maintains the domain, and some geographic information about the organization that the domain serves (such as the state or county). 726 domains were found in the CISA repository and 716 domains from the GSA repository. Because of how links were cleaned and aggregated, some .gov websites, particularly those corresponding to towns, cities, or counties, were not included in these repositories because they were subdomains of state root domains. We manually examined remaining websites ending with .gov and included any that correspond to a town, city, or county government, manually filling in relevant information about government type and geographic area.
Additionally, we identified some county and city websites using the .us domain that were not included in the GSA repository; because the .us domain is not strictly governmental, we manually reviewed all websites with the .us domain and included only those which corresponded to a county or city. 362 domains were added by manual review. In total, 1,848 websites were identified, which we call “government domains” or “government sources”. We are making the information about these domains in our dataset publicly available. [^18] https://bit.ly/FAccT24_gov_domains The distribution of domains across government types is shown in Table 2.
Government Domain Type Examples (geoprovenance) n
Federal whitehouse.gov (United States) 44
State az.gov (Arizona) 160
County cookcountyclerkil.gov (Cook County, Illinois) 697
Local cityofmaywood.com (Maywood, California) 806
Representative bishop.house.gov (Georgia, 2nd District) 137
Senator portman.senate.gov (Ohio) 3
Native nativefederation.org (Alaska) 1
Table 2. 1,848 government sources were identified from the data set of 4,556 domains (40.6%). Government sources were labeled for type and geoprovenance. We focus our analysis on state, county, and local government sources, as senator, representative, and Native sources show up minimally in organic results.
4.3. Quantifying Misjudgement
4.3.1. Domain level
Misjudgement can be studied to varying degrees of granularity. First, we examine misjudgement with respect to government domains and where they appear. To do so, we first identify how many government domains that appeared in organic results are from states in which we did not collect data; these are domains which are always mistargeted. Next, for domains originating in states for which we collected data, we calculate the proportion of times the domain appeared in the appropriate state and the proportion of times the domain was mistargeted (at the state-level). This gives insight into whether there are some locality-specific government domains that Google shows often across locations and are frequently mistargeted.
4.3.2. Aggregate Organic Results
We also study misjudgement by looking at all organic results in aggregate, and parsing differences in misjudgement by query categories, domain type, and state. Every organic result is labeled as to whether it is a government domain, and if so, we compare the domain’s geoprovenance (the geographic area it serves, such as a state, county, or city) to the search location. This gives us a more precise understanding of mistargeting than at the domain level, in which we only compare states. Organic results are labeled as either correct or mistargeted. Mistargeted results for county and city domains are also labeled as either in-state, if they serve a different location in the same state as the search location, or out-of-state, if they do not, which functions as an indicator for the severity of mistargeting. For each query category and state, as well as overall, we calculate the proportion of organic results composed of all non-federal government sources and the proportion of results composed of correctly targeted government results. Mistargeting rate is calculated by subtracting the proportion of correctly targeted from the total proportion, and dividing by the total proportion. We similarly calculate the proportion of results that are mistargeted but in-state, which gives insight into the nature of mistargeted results.
4.3.3. SERP Level
Lastly, we quantify misjudgement at the SERP level by calculating descriptive statistics about the number of government sources that appear per SERP, as well as descriptive statistics about the rank of correctly and incorrectly targeted government results. Since users heavily rely on ranking of search results (Pan et al., 2007), this informs the severity of misjudgement with respect to a user.
5. Results
5.1. RQ1: To what extent are government sources represented in the online environment of electoral information?
Of the 4,556 domains that appeared in organic results, 1,848 were identified as a government domain (40.6%). Overall, links belonging to these 1,848 government domains compose 39.97% of all organic results captured in the 3.45 million SERPs (31.3 million organic results) analyzed. Non-federal government domains compose 32.89% of all organic results. However, the extent to which government domains appear depends on the type of query, as well as the type of government source as shown in Table 3. In particular, government sources constitute 60.55% of organic results for Voting Logistics queries, with state government sources composing half of such appearances. In contrast, government sources are underrepresented among organic results for General Election Topics and Candidates and Issues queries, composing just 26.36% and 23.15% of organic results, respectively.
Voting Logistics General Election Topics Candidates and Issues All Queries
% Federal 4.85 9.1 8.44 7.08
% State 30.06 11.99 11.87 19.67
% County 22.91 4.75 2.66 11.86
% Local 2.62 0.52 0.18 1.31
% Total excl. Federal 55.7 17.26 14.71 32.89
% Total 60.55 26.36 23.15 39.97
Table 3. Government domains compose 39.97% of organic results overall, with state and county government domains constituting the majority of government domain appearances, while local government domains are minimally represented in organic results. The degree to which government domains appear depends largely on the query type: government sources compose 60% of organic results for Voting Logistics queries but only around a quarter for General Election Topics and Candidates and Issues.
Across all queries, local government sources are minimally represented among organic results. The prominence of state-level government sources compared to other government sources can be explained by the key roles that state-level actors play in elections: while the composition of election administrations vary by state, elections in each state are overseen by a chief election officer (such as a secretary of state, a governor, or a board), and their responsibilities include enforcing election laws, administering a statewide voter registration database, and other duties related to election logistics. [^19] https://www.ncsl.org/elections-and-campaigns/election-administration-at-state-and-local-levels However, their dominance is striking. While state government websites make up 3.5% of all websites (160 out of 4,556), they are shown in 30.06% of Voting Logistics results and in 19.67% of overall results.
On a SERP level, Voting Logistics queries had a median of 6 government results, with medians of 3 state-level and 2 county-level government sources, aligning with the aggregate proportions in Table3. In contrast, General Election Topics queries had a median of 2 government source per page and Candidates and Issues queries had a median of 1. The fact that the median of state-level websites per SERP is 3 for Voting Logistics queries naturally raises the question of whether all these three websites are targeted correctly to the search location.
5.2. RQ2: To what extent are state, county, and local government web domains correctly targeted to their respective constituents?
5.2.1. Domains
At the domain level, 174 of the 1,804 (9.6%) non-federal government sources represent a state for which we did not collect data; in other words, these sources were always mistargeted. This includes 71 local-level, 53 county-level and 48 state-level government domains, along with 2 senator websites. Appearances from these domains constitute 13.65% of appearances of non-federal government sources, or about 4.4% of all organic results. The remaining 1,630 non-federal government sources match, at the state level, the states for which we collected SERPs. 1,268 of these domains always show up in the appropriate state (although not necessarily in the appropriate county or locality, for county and local sources). While this is the majority of non-federal government domains (70.3%), only 4.96% of all organic appearances by non-federal government sources can be attributed to these domains.
Domain Type
State
Domain Serves
% Appearances
Correct State
% Appearances
Other States
Collected data
in state?
pa.gov State Pennsylvania 55.25 44.75 Yes
ca.gov State California 61.98 38.02 Yes
sos.state.tx.us State Texas 18.86 81.14 Yes
wi.gov State Wisconsin 26.30 73.69 Yes
wakegov.com County North Carolina 5.78 94.52 Yes
lavote.gov County California 25.79 74.21 Yes
ncsbe.gov State North Carolina 34.45 65.55 Yes
georgia.gov State Georgia 28.28 71.71 Yes
maryland.gov State Maryland 0 100.00 No
votebrevard.gov County Florida 5.83 94.17 Yes
Table 4. These 10 domains constitute 35.96% of all non-federal government organic results, which is 11.8% of all organic results. While collecting from more locations in certain states (such as Pennsylvania and California) increases the number of times these domains occur, all of these domains appear often in states for which they do not serve, indicating frequent mistargeting. maryland.gov is one of 174 non-federal sources that appear despite no data being collected for its respective state.
Table 4 shows the 10 non-federal government domains which make up the highest proportion of organic results, and together compose 35.96% of appearances for non-federal domains, or 11.8% of organic results overall. Because the number of locations for which data was collected varied per state, it is not appropriate to directly compare the frequency with which these domains occur to each other. However, it is clear that these domains suffer from high rates of mistargeting, with many appearing more often in other states than the states for which they serve. Notably, wakegov.com, a county government site for Wake County, North Carolina, and votebrevard.gov, the Supervisor of Elections site for Brevard County, Florida, are mistargeted over 94% of the time. Overall, 84 domains were mistargeted more often than they were correctly targeted, and they compose 44% of non-federal organic results (14.45% of all organic results). In contrast, 245 domains are correctly targeted more often than they are mistargeted, and they compose 41% of non-federal organic results (13.5% of all organic results).
Thus, while the majority of non-federal government sources are always correctly targeted at the state-level, these sources make up only a minority of all organic results; a relatively small subset of domains are mistargeted more often than they are correctly targeted, and these mistargeted results show frequently among organic results.
5.2.2. Misjudgement Levels by Query Category and State
Table 5 shows the percent of all organic results that are correctly targeted to the appropriate locality, and Table 6 shows the mistargeting rates for each state, county, and local government sources for different query categories. 71% of all appearances of locality-specific government websites in organic results were mistargeted, with the proportion of mistargeting highest for Candidates and Issues queries. Voting Logistics queries, which see the highest proportion of government results by far, have the lowest mistargeting rate of the 3 query categories, albeit still 69.4%.
Voting Logistics General Election Topics Candidates and Issues All Queries
% State Correct 14.5 4.1 3.0 8.2
% County Correct 1.89 0.42 0.11 0.96
% Local Correct 0.63 0.13 0.03 0.31
% Total Correct 17.02 4.65 3.14 9.47
Table 5. Percent of all organic results that are correctly targeted to the appropriate locality for state, county, and local sources. While 32.89% of all organic results can be attributed to locality-specific government sources, only 9.47% of organic results are composed of correctly targeted government sources.
Voting Logistics General Election Topics Candidates and Issues All Queries
% State Mistargeted 51.75 65.8 74.73 58.29
% County Mistargeted 91.73 91.12 95.7 91.94
% Local Mistargeted 75.92 75.06 81.79 76.08
% Total Mistargeted 69.4 73.05 78.61 71.18
Table 6. Proportion of government results that were mistargeted; the denominator here is the total appearances of a source type, rather than all organic results. Overall, 71% of locality-specific government organic results were mistargeted, with the mistargeting rate being highest for county government sources at nearly 92%.
Across all queries, county-level government sources suffer from the highest levels of mistargeting, with 91.9% of these results mistargted. State-level government domains, which appear the most among organic results, have the lowest mistargeting rate of the locality-specific sources at about 58.3%. Mistargeting of state-level government sources is lowest for Voting Logistics queries at 51.75%, notable since state-level sources are the most common government source in this query category. Lastly, local government sources suffer from a mistargeting rate of 76%, unfortunate considering the minimal appearances of local government sources in organic results.
There is some variation in the extent to which government domains appear in different states, as well as mistargeting rates by state, as visualized in Figure 2. Studying this variation is left to future work.
Figure 2. Variation in the extent to which government sources appear in different states. much of the mistargeting for local government sources (and county government sources, to a lesser extent) can be attributed to other government sources within the respective state, indicating a lack of precision with targeting locality-specific government domains to the appropriate audience when relying only on geocoordinates to represent user location.
5.2.3. Mistargeting Severity
While mistargeting rates are high, particularly for county and local government domains, it is useful to distinguish between error rate and error severity. Error severity can be assessed in two ways: 1) the rank of mistargeted organic results, 2) how close mistargeted results are to the correct location.
With respect to rank, while only 28.8% of non-federal government sources are correctly targeted, correctly targeted results generally appear in higher ranks among organic results than mistargeted results (Figure 3). The median rank of a correctly-targeted government result was 3, while the median rank of a mistargeted government result was 7. Moreover, for the 43.7% of SERPs that contained both correctly-targeted and mistargeted government results, the first correctly-targeted result appeared above the first mistargeted result 83.9% of the time. A mistargeted government result appeared in the first organic position for only 3.3% of SERPs analyzed, while a correctly-targeted government result appeared in the first organic position for 19.3% of SERPs. Given users tend to focus on search results in the top positions (Pan et al., 2007), these findings suggest that while mistargeting rates are high, their severity, within the context of impact on a user, may be low. Additionally, these findings may indicate that the search algorithm prioritizes non-federal government sources in ranked results only when there is a high likelihood that the source geoprovenance matches the user’s locations; otherwise, non-federal government sources appear in lower ranked positions, and other content (such as a news or non-profit organization) is given the top ranks.
Figure 3. While only 28.8% of non-federal government domains are correctly targeted, these results usually appear in higher ranks among organic results. On the 43.7% of SERPs that have both correctly-targeted and mistargeted results, the correctly targeted result appeared above the mistargeted result 83.9% of the time.
With respect to the nature of mistargeted results, a high percentage of mistargeting for local government sources can be attributed to local government sources for places in the same state as the search location. While the overall mistargeting rate for local government sources is 76%, 64% of local government results are for other localities in the correct state. Figure 2 visualizes this phenomenon; in most states, nearly all of the mistargeted local government results are from other local government sources in the respective state. A similar phenomenon with targeting county-level government sources is observed, though to a lesser degree and with more variation between states. These findings imply a lack of precision with targeting locality-specific government sources to the appropriate audience when using general queries that do not specify location information. While a similar lack of precision has been observed with geotargeting of Ads on Google (Bandy and Hecht, 2021), the consequences of this lack of precision is greater in the context of “low information voters” [^20] https://en.wikipedia.org/wiki/Low_information_voter searching for election information as it relates to their locale, given that other county or locality information is likely not be applicable to them.
Lastly, while we find no correlation between the number of local governments and the percent that local government sources compose in organic results, we find a moderate, positive correlation between the number of local governments in a state and the mistargeting rate for local government sources (r=0.36), and for the correlation between number of local governments and the rate of in-state mistargeting (r=0.37), suggesting to some degree that a greater number of local governments is associated with less precise targeting of local government sources in organic search.
5.3. Presence of Election Information Boxes
Between Oct 18-24, only 39 queries showed an electoral box (across most locations), though this number jumped to 73 on Oct 25 (two weeks before election) and went up to 83 on election day. Notably, on election day, all locations had at least one election box. Once the elections were over, the number of queries with an election box went back in the 30s, but remained consistent across most locations.
These results indicate that the presence of election information boxes is not location-specific, but depends on the date and query phrase. Only 86 out of 161 query phrases (53%) had at least one box on at least one date. Furthermore, 66% of these queries (57 out of 86) belonged to the list of queries we collected from Google Trends. These queries made up 77% of all queries from that source (57 out of 74). This suggests that Google might be showing election information boxes for popular query phrases of a certain type. However, a more detailed analysis (with a more expansive set of queries) might be needed to understand what triggers these boxes.
6. Discussion and Future Work
6.1. How Algorithmic Misjudgement Shapes the Electoral Information Environment
The results of our first research question reveal that about 40.6% of all websites are government sources and that combined they contribute 39.97% of organic results. This might look like an excellent outcome, given the almost perfect matching between the proportion of domains and the proportion of their occurrences in the search results. However, our deeper analysis of the results reveals several issues that are hidden, when one only considers proportions.
For one, the top ten non-federal government sources show up disproportionally often, contributing to 11.8% of all organic results (see Table 4). As a reference point, all 1,848 non-federal sources combined make up only 32.89% of all results (i.e., about 1/3 of their combined results are accounted for by these ten domains). This phenomenon is reminiscent of source concentration and mainstreaming effects found in previous work (Trielli and Diakopoulos, 2022, 2019) for news sources. While it might not be harmful to an individual to read national news from New York Times more than from Baltimore Sun (although, collectively, source concentration has led to the so-called ’news desert’ phenomenon (Abernathy, 2018)), for elections, all out-of-state web domains might be irrelevant to a voter, when it comes to their voting information. Our results make it clear that Google Search is treating electoral information needs as any other type of information and is populating the organic search results with popular websites, instead of locally-relevant websites.
Further, while correctly-targeted government sources are usually ranked higher than those that are mistargeted (Figure 3), indicating that Google is capable of getting things right (19.3% of SERPs have a correctly targeted government website at the first position, mostly a state government), 71.19% of government website occurrences were mistargeted. Overall, our analysis indicates a lack of consistency in the search algorithm’s treatment of locality-specific government sources, leaving most government domains, especially those of local governments, appearing minimally, inconsistently, or missing entirely from election-related search results. We consider the case of mistargeted websites and that of the absence of locality-specific websites as instances of algorithmic misjudgement that need to be corrected.
Google Search is aware of the importance of authoritative information about elections, as indicated by the consistent inclusion of “election information panels” in prominent SERP positions across locations. Google should consider that showing locality-specific web domains in the organic results might contribute to strengthening the civic infrastructure of the election by increasing the visibility of information from elected officials.
6.2. Limitations
There are limitations to our study. First, we focus exclusively on government web domains in search results, but in reality there are other types of information that users encounter in search results, including results from other types of sources and curated information panels. Additionally, our audit was conducted when there were around 10 organic results per SERP, while since December 2022 (after our data collection), Google has moved to “continous scrolling”, which loads up to 60 results at once. [^21] https://www.theverge.com/2022/12/6/23495973/google-search-desktop-continuous-scrolling Second, we rely on geocoordinates to serve as a proxy for locations of real users and do not study how including location information in the search query (eg. “voting near me Boston, MA”) impacts the composition and mistargeting of government search results. While the queries come from Google Trends and real users, we do not study how users actually interact with government sources in search results. Moreover, differences in popularity of search terms between those that come from Google Trends and those from real users may impact the composition of search results (i.e. more popular queries may have better results).Third, we analyze data from 786 locations across 25 states, a size which is sufficient for statistical analysis but does not capture all of the United States. Finally, this study only focuses on the United States, given that electoral laws and election authorities are different state by state, making this kind of analysis meaningful. In other countries, where there is a single election authority, the needs and assumptions might be different.
6.3. Future work
One challenge that we encountered when analyzing our results was finding accurate information about government websites (as discussed in Methods). However, government websites were only 40.6% of the entire dataset of websites. The rest was composed of news organizations, non-profit civic organizations, companies, universities, individual blogs, and more. There are no high-quality and freely available databases that provide detailed information on web domains (their nature and what audiences they serve). Nuanced audits of search engine results need such information; therefore, methods to create and maintain such databases need to be developed. (We tried to use some news related curated lists for other studies, such as (Fischer et al., 2020), but their overlap with our dataset was not high.)
A scraping audit is useful for controlling the inputs to the algorithmic system, so that we can systematically compare results, but it does not involve real voters with real information needs, that is, it lacks ecological validity. Currently, there are efforts underway like the National Internet Observatory [^22] https://nationalinternetobservatory.org/ that will enable researchers to carry out crowdsourced audits, similar to what was used in (Robertson et al., 2023). Future audits can then address both issues that are central to our audit: what queries do users formulate (and reformulate) to find information about elections and with what results do they engage? For example, do they click on Google’s election boxes, do they click on their own local/county/state websites, or do they visit non-governmental sources? Additionally, such audits might allow us to find out if Google Search is bypassed entirely and users go directly to their trusted sources.
6.4. Ethical Considerations
Our study does not include any personal data, as Google Search result pages are public data. Google handles circa 8.5 billion of searches per day, [^23] https://blog.hubspot.com/marketing/google-search-statistics so our scraping audit of 140 K daily searches provided a minimal burden. Finally, a US Federal Court has ruled that audits such as ours are legal and do not violate the Computer Fraud and Abuse Act. [^24] https://www.aclu.org/press-releases/federal-court-rules-big-data-discrimination-studies-do-not-violate-federal-anti
Acknowledgements.
We are grateful to members of our lab and funding from XXX.